What is the final temperature of the water if 35.1 kJ of energy are added to 100.0 g of ice at -4.5°C?


1
Expert's answer
2020-12-01T09:35:42-0500

Quantity of Heat= mici(Tfi-Tii) + mLf + mwcw(Tfw-Tiw)

                    35.1= 0.1×2.04× (0- -4.5) + 0.1×335 +0.1×4.184(Tfw-0)

                       35.1= 0.918 + 33.5 +0.4184 Tfw

           35.1-34.418= 0.4184 Tfw

             0.682/0.4184= Tfw

             Tfw= 1.63 °C